(4) An unknown gas occupies a volume of 4.75 L at 1500 K and 5.00 atm. If its mass is 5.45 g, what is the molar mass of the gas? Hint: molar mass = dRT/P.



1
Expert's answer
2021-05-06T07:39:38-0400

We have given the volume V = 4.75 L

Pressure = 5 atm

Mass in gram = 5.45 gram


R = 0.02821 atmL/molK

Temperature T =1227°C=1227+273=1500 K

From ideal gas equation PV=nRT

"5\times 4.75=n\times 0.0821\times 1500"


"5\u00d74.75=n\u00d70.0821\u00d71500"


n=0.1928


So molar mass "\frac{mass\ in\ gram}{n}=\frac{5.45}{0.1928}=28.25g/mol"